You are choosing between two projects. The cash flows for the projects are given in the following table ($ million):

Project

Year 0

Year 1

Year 2

Year 3

Year 4

A

−$50

$25

$21

$18

$16

B

n−$98

$19

$38

$52

$59

a. What are the IRRs of the two projects?

b. If your discount rate is 4.9%, what are the NPVs of the two projects?

c. Why do IRR and NPV rank the two projects differently?
Answer #1

- fx =IRR(B3:37) 1 Project A 2 year Cash flows pv@4.9% Present value (50.00) 1.0000 $ -50.00 25.00 0.9533 $ 23.83 21.00 0.908

B8 fx IRR(B3:37) pv@4.9% 1 Project A 2 year 30 41 52 63 Cash flows -50 25 21 18 =C3/1.049 =C4/1.049 =C5/1.049 =C6/1.049 NPV P